1. Solve: (x+4) = 3 a) -1, -7 b) - 4+ 13 C) 43 13 d) - 2 + V3 e) -13,5
Answer #1

san (x+494 = 3 ie (x+4)²- 320 (x+4)2-(13)2 = 0 (22+4-13) (X+ 4+ v3) =0 x+4-3=0 a+4+3 =0 x = -4 +13 76 =-4-13 ) iz x=-42 ßß ol